# Centre du Parc, Rue Marconi 19, CH-1920 Martigny, Switzerland

### v1

@@ 5.1 Obstacle Avoidance (PNGRID)
Table 1 presents a comparative analysis of system performance across baseline approaches, based on averages from 100 trials with random start and target points. All Proj-MPPI, NF-PoE-MPPI and TT-PoE-MPPI demonstrate performance improvements over the MPPI method, indicated by negative values, consistent with prior findings (Power and Berenson, 2023; Sacks and Boots, 2023). Notably, the TT-PoE-MPPI method consistently outperforms the other methods, particularly in terms of success rate within the medium to low sampling regime. However, as the number of samples increases, both NFs-PoE-MPPI and TT-PoE-MPPI exhibit improved performance in terms of step efficiency and cost values. In contrast, Proj-MPPI’s performance aligns more closely with MPPI under high-sampling conditions than in the lower and medium sampling regimes. This trend highlights the competitive advantage of the TT-PoE-MPPI and NFs-PoE-MPPI approaches across varying sampling densities.
5.2 Moving on manifolds (e.g., surface cleaning or painting)
The proposed method is also applied to a task where the robot’s end-effector must traverse a manifold (e.g., a spherical surface or a sinusoidal path, as illustrated in Fig. 10). Such tasks are prevalent in robotics. For instance, when a robot needs to clean or paint a surface or follow a specific path. In this setup, we control the robot’s end-effector position in the task space, ensuring it stays within a designated shaded area. However, the feasible region for this task is substantially restricted relative to the entire workspace, making it challenging to sample feasible actions. We employed two different manifolds: one between two spheres and another between two sinusoidal waves (Fig. 10). The evaluation results, shown in Table 1, reveal that the proposed method significantly improves the task success rate, especially in low-sampling scenarios. Moreover, our method achieves faster convergence with lower associated costs. As shown in the table, Proj-MPPI exhibits behavior similar to MPPI in terms of success rate (and even lower in the sphere case) and number of steps, while achieving significantly lower cost values. This outcome is likely due to the projection step, which ensures that nearly all samples remain within the designated area, leading to reduced costs. However, this projection mechanism appears to make the system more inclined to stay within its current location rather than actively explore new tasks, which may contribute to its lower success rate on the sphere manifold compared to MPPI. In contrast, NFs-PoE-MPPI shows markedly lower performance than TT-PoE-MPPI, for reasons that are discussed in detail in Section 6.
@@
succ. succ. log(#$\#\text{steps}_{n}$) log($\text{cost}_{n}$) succ. log(#$\#\text{steps}_{n}$) log($\text{cost}_{n}$) succ. log(#$\#\text{steps}_{n}$) log($\text{cost}_{n}$)
[…]
@@ 5.3 Non-prehensile Manipulation (Planar-pushing)
The second task that we consider is a pushing task (Mason, 1986; Xue et al., 2023), a type of non-prehensile, contact-rich manipulation. These tasks are inherently challenging as they require the robot to handle a range of contact models, with hybrid dynamics that complicate the use of gradient-based methods. This hybrid nature arises from various factors, including the geometry of objects, the contact modes, as well as the task level. In this work, we focus specifically on the challenges related to geometry and contact mode, while task-level hybridization has been more extensively explored within task and motion planning (TAMP) frameworks (Holladay et al., 2023).
Previous studies have attempted to address the hybrid nature of contact through unified contact modes (Anitescu, 2006; Manchester and Kuindersma, 2020; Posa et al., 2014) or by employing mixed-integer programming (MIP) to jointly optimize hybrid and continuous variables. However, these approaches struggle with geometric hybridness, and solving complex non-linear, non-convex problems with MIP can be computationally intensive. SMPC offers an effective alternative (Pezzato et al., 2023), as it avoids reliance on gradients. Furthermore, as demonstrated by Pang et al. (2023), the inherent ability of sampling-based methods to smooth functions proves advantageous in these hybrid and contact-rich scenarios.
Using the method developed in this work, the optimality expert can focus on pushing an object toward a goal while adapting to environmental changes, such as the introduction of obstacles that were not present in the training data. The feasibility distribution is designed to model the efficiency of the actions taken. As mentioned earlier, certain actions are inconsequential if they lead the robot away from the object or otherwise fail to move the object meaningfully. In an offline phase, we gather data on various state-action combinations, labeling them as efficient if the robot moves toward the object or if the action causes the object to relocate. Specifically, $p(\bm{x}_{t},\bm{u}_{t})$ in (13) returns 1 if the action $\bm{u}_{t}$ in state $\bm{x}_{t}$ is effective and 0 otherwise. This approach is similar to embedding a cost function that encourages the robot to approach the object. However, we demonstrate that our method is more efficient than relying on a penalty term within the cost function.
In this experimental setup, we employed three objects: a cylinder, a box, and a mustard bottle sourced from the YCB dataset (Calli et al., 2017). Additionally, for the cylinder scenario, we introduced an obstacle into the environment to assess the efficacy of the methods in navigating around obstacles without collisions. It is important to note that the feasibility distribution is unaware of the obstacle in the environment (not considered in the training), as addressing this situation falls within the responsibility of the optimality distribution.
@@
The proposed method was further applied to the real robot for pushing a mustard bottle using only 32 samples which is very low compared to the commonly used number of samples in other SMPC methods. We chose this low number deliberately to show the capability of our method in increasing the efficiency of samples. Selected frames from this experiment are illustrated in Fig. 9. The rationale behind selecting the mustard bottle lies in its complex shape compared to other objects, thereby serving as a demonstrative example that the shapes of the objects involved do not constrain our method. The objective in this scenario is to guide the object back to the goal location (transparent shape in Fig. 9), while the object undergoes a 90-degree rotation. The receding horizon controller employed allows the system to effectively handle external disturbances, as demonstrated in the accompanying video. Disturbances were introduced to the system between Fig. 9-c and 9-d. Notably, the planner adeptly manages this situation, guiding the object towards the target.
@@
Figure 10: Environmental setups for movements on manifold: (a) sphere and (b) sinusoidal wave.
Task #samples MPPI NFs-PoE-MPPI TT-PoE-MPPI
succ. rate succ. rate log(#$\#\text{steps}_{n}$) log($\text{cost}_{n}$) succ. rate log(#$\#\text{steps}_{n}$) log($\text{cost}_{n}$)
[…]
@@ 5.4 Online Obstacle Avoidance
One significant advantage of using TT-SVD to learn the TT distribution is its capability to process the entire dataset in milliseconds if all data points are available. This is especially beneficial when the robot leverages point cloud data online to identify free space in the environment. This rapid processing allows the system to adapt its distribution quickly if new environmental changes are detected, enabling it to sample updated, obstacle-free trajectories efficiently. To evaluate this capability, we designed a task in which a point-mass robot moves from the left to the right side of the environment. As it progresses, new obstacles appear, shown as lighter regions in Fig. 1-b. Table 3 presents the results of this experiment, demonstrating that the proposed method consistently outperforms the MPPI method across all measured criteria. Notably, the performance gap between the two methods widens as the number of samples increases.
@@
Table 3: Comparison of TT-PoE-MPPI with MPPI in an online obstacle avoidance task across 100 random obstacle locations. Values are normalized to the MPPI method. Thus, steps and cost values for MPPI are omitted. Negative values indicate superior performance, with bold formatting highlighting the best results.
#samples MPPI TT-PoE-MPPI
succ. rate succ. rate log(# $\text{steps}_{n}$) log($\text{cost}_{n}$)
[…]

### v2

@@ 5.1 Obstacle Avoidance (PNGRID)
Table 1 presents a comparative analysis of system performance across baseline approaches, based on averages from 100 trials with random start and target points. All Proj-MPPI, NFs-PoE-MPPI and TT-PoE-MPPI demonstrate performance improvements over the MPPI method, indicated by negative values, consistent with prior findings (Power and Berenson, 2023; Sacks and Boots, 2023). Notably, the TT-PoE-MPPI method consistently outperforms the other methods, particularly in terms of success rate within the medium to low sampling regime. However, as the number of samples increases, both NFs-PoE-MPPI and TT-PoE-MPPI exhibit improved performance in terms of step efficiency and cost values. In contrast, Proj-MPPI’s performance aligns more closely with MPPI under high-sampling conditions than in the lower and medium sampling regimes. This trend highlights the competitive advantage of the TT-PoE-MPPI and NFs-PoE-MPPI approaches across varying sampling densities.
@@
succ. succ. log($\text{steps}_{n}$) log($\text{cost}_{n}$) succ. log($\text{steps}_{n}$) log($\text{cost}_{n}$) succ. log($\text{steps}_{n}$) log($\text{cost}_{n}$)
  PNGRID 16 $46\%$ $70\%$ 0.07-0.07 0.19-0.19 $74\%$ 0.03-0.03 0.04-0.04 $\bm{96\%}$ $\bm{-0.81}$ $\bm{-0.35}$
64 $81\%$ $92\%$ 0.06-0.06 0.09-0.09 $90\%$ 0.10-0.10 0.16-0.16 $\bm{100\%}$ $\bm{-0.69}$ $\bm{-0.90}$
512 $93\%$ $96\%$ 0.06-0.06 0.09-0.09 $97\%$ 0.13-0.13 0.16-0.16 $\bm{100\%}$ $\bm{-0.65}$ $\bm{-0.78}$
  Sphere. Man. 16 $\% 7373\%$ 0.11-0.11 0.09-0.09 $75\%$ $0.01$ 0.57-0.57 $\bm{100\%}$ $\bm{-0.32}$ $\bm{-1.08}$
64 $93\%$ $77\%$ $0.03$ 1.57-1.57 $87\%$ $0.20$ 0.68-0.68 $\bm{100\%}$ $\bm{-0.39}$ $\bm{-4.38}$
512 $\bm{100\%}$ $78\%$ $0.05$ 1.89-1.89 $\bm{100\%}$ 0.26-0.26 1.97-1.97 $\bm{100\%}$ $\bm{-0.89}$ $\bm{-3.08}$
  Sin. Man. 16 $31\%$ $38\%$ 0.09-0.09 1.08-1.08 $38\%$ 0.15-0.15 0.08-0.08 $\bm{100\%}$ $\bm{-1.58}$ $\bm{-1.86}$
64 $\% 4444\%$ 0.09-0.09 0.87-0.87 $50\%$ 0.41-0.41 0.40-0.40 $\bm{100\%}$ $\bm{-2.37}$ $\bm{-2.03}$
512 $55\%$ $68\%$ 0.30-0.30 0.68-0.68 $65\%$ 0.52-0.52 0.14-0.14 $\bm{100\%}$ $\bm{-1.92}$ $\bm{-2.40}$
5.2 Tracking Tube (e.g., surface cleaning or painting)
The proposed method is also applied to a task where the robot’s end-effector must traverse a restricted volume (e.g., a spherical or a sinusoidal manifold with some margins, as illustrated in Fig. 8). Such tasks are prevalent in robotics. For instance, when a robot needs to clean or paint a surface or follow a specific path. In this setup, we control the robot’s end-effector position in the task space, ensuring it stays within a designated shaded area. However, the feasible region for this task is substantially restricted relative to the entire workspace, making it challenging to sample feasible actions. We employed two different volumes: one between two spheres and another between two sinusoidal waves (Fig. 8). The evaluation results, shown in Table 1, reveal that the proposed method significantly improves the task success rate, especially in low-sampling scenarios. Moreover, our method achieves faster convergence with lower associated costs. As shown in the table, Proj-MPPI exhibits behavior similar to MPPI in terms of success rate (and even lower in the sphere case) and number of steps, while achieving significantly lower cost values. This outcome is likely due to the projection step, which ensures that nearly all samples remain within the designated area, leading to reduced costs. However, this projection mechanism appears to make the system more inclined to stay within its current location rather than actively exploring new solutions, which may contribute to its lower success rate on the sphere manifold compared to MPPI. In contrast, NFs-PoE-MPPI shows markedly lower performance than TT-PoE-MPPI, for reasons that are discussed in detail in Section 6.
In this experiment, our objective was to evaluate the performance of our systems on equality constraints, which in this case correspond to a manifold. However, we acknowledge that the TT method, which is a volumetric representation, cannot directly enforce strict equality constraints. To address this limitation, we evaluated the performance of our approach on a relaxed version of the problem by introducing margin boundaries around the manifold, effectively creating a ”soft” constraint region for analysis. We then evaluate how the pipeline behaves as this margin is progressively reduced, and benchmark it against an intrinsic‐sampling approach (i.e., Riemannian log/exp maps), which guarantees constraint satisfaction intrinsically by sampling in the tangent space first, then mapping to the task space. For certain manifolds, like a sphere, these mappings can be computed analytically (see Appendix for details).
Table 2 presents the results of this comparison. We evaluate our proposed method against the Riemannian log/exp maps approach (specifically for a spherical manifold) and a vanilla-MPPI method. As shown in Table 2, when the margin exceeds the TT discretization (2 cm), our TT‑PoE‑MPPI method outperforms both standard MPPI and the Riemannian mapping approach. We attribute this to the fact that intrinsic sampling in the tangent space incurs inefficiencies, particularly when shifting between time steps and averaging, whereas the TT representation captures the volumetric structure more globally. As the margin shrinks, however, the intrinsic method ultimately yields the best performance, although our method remains superior to vanilla MPPI.
These results suggest that TT‑PoE‑MPPI can accelerate the dynamic sampler under equality constraints, but that pure intrinsic sampling may be preferable for very tight margins. Moreover, TT‑PoE can be combined with intrinsic approaches: once an intrinsic transform renders the problem unconstrained, additional constraints (e.g., joint limits or other inequalities) can still be enforced by defining the TT experts in the transformed space. Note tha the optimality distributions for the intrinsic sampler and for other methods inhabit different spaces in this experiment. Hence, as the margin shrinks to zero, the unconstrained intrinsic sampler can serve as the ground-truth solution to the constrained problem. Furthermore, our product‑of‑experts framework consistently improves sampling efficiency without imposing additional structural assumptions, such as that the constraint space forms a Riemannian manifold or that the constraints are affine in the action variables, as required by Carius et al. (2021).
Table 2: Comparison of different methods on sphere tube-tracking task across varying margin sizes. The intrinsic MPPI method (MPPI solved directly on the spherical manifold using log/exp maps) reformulates the problem as unconstrained and, in the limit as the margin $\to 0$, serves as the ground truth for the other baselines. Values are normalized to the MPPI method. Hence, the number of steps and cost values for the MPPI method are not provided. Negative values indicate superior performance compared to the MPPI method, with bold formatting indicating superior results.
Margin size MPPI $T$-PoE-MPPI Intrinsic MPPI
succ. rate succ. rate log($\text{steps}_{n}$) log($\text{cost}_{n}$) succ. rate log($\text{steps}_{n}$) log($\text{cost}_{n}$)
  5 [cm] $98\%$} $\bm{100\%}$ $\bm{-0.98}$ $\bm{-2.94}$ $\bm{100\%}$ 0.72-0.72 2.55-2.55
  3 [cm] $91\%$} $\bm{100\%}$ $\bm{-1.06}$ $\bm{-4.03}$ $\bm{100\%}$ 0.81-0.81 3.78-3.78
  1 [cm] $78\%$} $92\%$ 0.02-0.02 4.91-4.91 $\bm{100\%}$ $\bm{-0.37}$ $\bm{-5.73}$
Figure 8: Different environmental setups: (a) sphere tube tracking, (b) sinusoidal wave tube tracking, and (c) Online obstacle-avoidance. In (c), the obstacles become more apparent as the robot approaches them; lighter colors indicate obstacles that appear later in the task.
@@ 5.3 Non-prehensile Manipulation (Planar-pushing)
An other task that we consider is a pushing task (Mason, 1986; Xue et al., 2023), a type of non-prehensile, contact-rich manipulation. These tasks are inherently challenging as they require the robot to handle a range of contact models, with hybrid dynamics that complicate the use of gradient-based methods. This hybrid nature arises from various factors, including the geometry of objects, the contact modes, as well as the task level. In this work, we focus specifically on the challenges related to geometry and contact mode, while task-level hybridization has been more extensively explored within task and motion planning (TAMP) frameworks (Holladay et al., 2023).
Previous studies have attempted to address the hybrid nature of contact through unified contact modes (Anitescu, 2006; Manchester and Kuindersma, 2020; Posa et al., 2014; Yunt and Glocker, 2006) or by employing mixed-integer programming (MIP) to jointly optimize hybrid and continuous variables (Aceituno and Rodriguez, 2022). However, these approaches struggle with geometric hybridness, and solving complex non-linear, non-convex problems with MIP can be computationally intensive. SMPC offers an effective alternative (Pezzato et al., 2023), as it avoids reliance on gradients. Furthermore, as demonstrated by Pang et al. (2023), the inherent ability of sampling-based methods to smooth functions proves advantageous in these hybrid and contact-rich scenarios.
Using the method developed in this work, the optimality expert can focus on pushing an object toward a goal while adapting to environmental changes, such as the introduction of obstacles that were not present in the training data. The feasibility distribution is designed to model the efficiency of the actions taken. As mentioned earlier, certain actions are inconsequential if they lead the robot away from the object or otherwise fail to move the object meaningfully. In an offline phase, we gather data on various state-action combinations, labeling them as efficient if the robot moves toward the object or if the action causes the object to relocate. Specifically, $p_{\text{feas}}(\bm{x}_{t},\bm{u}_{t})$ in (13) returns 1 if the action $\bm{u}_{t}$ in state $\bm{x}_{t}$ is effective and 0 otherwise. This approach is similar to embedding a cost function that encourages the robot to approach the object. However, we demonstrate that our method is more efficient than relying on a penalty term within the cost function.
In this experimental setup, we employed three objects: a cylinder (Fig. 9-a), a box, and a mustard bottle sourced from the YCB dataset (Calli et al., 2017) (Fig. 9-b). Additionally, for the cylinder scenario, we introduced an obstacle into the environment to assess the efficacy of the methods in navigating around obstacles. It is important to note that the feasibility distribution is unaware of the obstacle in the environment (not considered in the training), as addressing this situation falls within the responsibility of the optimality distribution.
@@
Task #samples MPPI NFs-PoE-MPPI $T$-PoE-MPPI
succ. rate succ. rate log($\text{steps}_{n}$) log($\text{cost}_{n}$) succ. rate log($\text{steps}_{n}$) log($\text{cost}_{n}$)
  Push. Cir. 16 $89\%$ $\bm{96\%}$ 0.19-0.19 0.22-0.22 $95\%$ $\bm{-0.35}$ $\bm{-0.41}$
64 $6\% 99$$9\%$ 0.27-0.27 0.30-0.30 $\bm{100\%}$ $\bm{-0.38}$ $\bm{-0.45}$
512 $97\%$ $\bm{100\%}$ 0.24-0.24 0.27-0.27 $\bm{100\%}$ $\bm{-0.38}$ $\bm{-0.45}$
  Push. Box 16 $\bm{100\%}$ $\bm{100\%}$ $0.29$ $0.28$ $\bm{100\%}$ $\bm{-0.17}$ $\bm{-0.21}$
64 $\bm{100\%}$ $\bm{100\%}$ 0.06-0.06 0.08-0.08 $\bm{100\%}$ $\bm{-0.25}$ $\bm{-0.29}$
512 $\bm{100\%}$ $\bm{100\%}$ $0.25$ $0.23$ $\bm{100\%}$ $\bm{-0.01}$ $\bm{-0.03}$
  Push. Mus. 16 $70\%$ $73\%$ $0.12$ $0.11$ $\bm{89\%}$ $\bm{-0.21}$ $\bm{-0.25}$
64 $75\%$ $90\%$ 0.01-0.01 0.02-0.02 $\bm{91\%}$ $\bm{-0.33}$ $\bm{-0.34}$
512 $75\%$ $90\%$ $0.01$ 0.01-0.01 $\bm{91\%}$ $\bm{-0.26}$ $\bm{-0.28}$
The proposed method was further applied to the real robot for pushing a mustard bottle using only 32 samples which is very low compared to the commonly used number of samples in other SMPC methods. We chose this low number deliberately to show the capability of our method in increasing the efficiency of samples. Selected frames from this experiment are illustrated in Fig. 10. The rationale behind selecting the mustard bottle lies in its complex shape compared to other objects, thereby serving as a demonstrative example that the shapes of the objects involved do not constrain our method. The objective in this scenario is to guide the object back to the goal location (transparent shape in Fig. 10), while the object undergoes a 90-degree rotation. The receding horizon controller employed allows the system to effectively handle external disturbances, as demonstrated in the accompanying video. Disturbances were introduced to the system between Fig. 10-c and 10-d. Notably, the planner adeptly manages this situation, guiding the object towards the target.
@@
Table 4: Comparison of $T$-PoE-MPPI with MPPI in an online obstacle avoidance task across 100 random obstacle locations. Values are normalized to the MPPI method. Thus, steps and cost values for MPPI are omitted. Negative values indicate superior performance, with bold formatting highlighting the best results.
#samples MPPI $T$-PoE-MPPI
succ. rate succ. rate log(# $\text{steps}_{n}$) log($\text{cost}_{n}$)
  16 $73\%$ $\bm{84\%}$ $\bm{-0.09}$ $\bm{-0.13}$
64 $73\%$ $\bm{86\%}$ $\bm{-0.31}$ $\bm{-0.44}$
512 $69\%$ $\bm{89\%}$ $\bm{-0.46}$ $\bm{-0.66}$
@@ 5.4 Online Obstacle Avoidance
One significant advantage of using TT-SVD to learn the TT distribution is its capability to process the entire dataset in milliseconds if all data points are available. This is especially beneficial when the robot leverages point cloud data online to identify free space in the environment. This rapid processing allows the system to adapt its distribution quickly if new environmental changes are detected, enabling it to sample updated, obstacle-free trajectories efficiently. To evaluate this capability, we designed a task in which a point-mass robot moves from the left to the right side of the environment. As it progresses, new obstacles appear, shown as lighter regions in Fig. 8-c. Table 4 presents the results of this experiment, demonstrating that the proposed method consistently outperforms the MPPI method across all measured criteria. Notably, the performance gap between the two methods widens as the number of samples increases.
5.5 Higher Dimensions and Alternative Optimality Distributions
Previous experiments addressed low-dimensional problems using TT‑SVD and benchmarked against vanilla MPPI. Here, we demonstrate that our method scales to higher-dimensional settings and can be combined with alternative optimality distributions. We evaluate our method on a task in which a Franka Emika robot must traverse an environment with whole‑body obstacle avoidance, where obstacle locations vary across scenes (as illustrated Fig. 1-b).
Since TT‑SVD becomes impractical for the large tensors required, we instead employ TT‑cross with a fixed maximum rank $r_{\max}=20$. This low‑rank approximation reconstructs the full tensor from a subset of entries. Our approach can be applied in two ways:
1. Augmented State Conditioning: Augment the state with all obstacle positions and, at inference time, condition on the current robot state $\bm{x}_{t}$ and obstacle locations $\bm{x}^{o}_{1:N_{o}}$. We define the joint feasibility distribution as
$p^{f}\bigl(\bm{x}^{o}_{1:N_{o}},\bm{x}_{t},\bm{u}_{t}\bigr)=\bigl(1-\exp(-\mathrm{dist}(\bm{x}_{t})/0.5)\bigr)\\ \bigl(1-\exp(-\mathrm{dist}(\bm{x}_{t+1})/0.5)\bigr),$ (25)
where $\mathrm{dist}(\cdot)$ is the minimum distance between the robot body and obstacles computed via the Robot Distance Field (RDF) in (Li et al., 2024). By using a smooth, distance‑based likelihood rather than a binary collision indicator, we favor configurations that keep the robot at a safe distance and produce a lower‑rank function amenable to TT‑cross learning.
2. Additive Feasibility Mixture: Learn a separate feasibility tensor $\mathcal{P}^{f_{i}}$ for only one obstacle, then combine them as
$\mathcal{P}^{f}\;=\;\sum_{i=1}^{N_{o}}\,\mathcal{P}^{f_{i}}.$
The additive mixture remains smoother with a lower rank representation, which is more amenable to the TT-cross method. We denote this variant as TT‑PoE‑MPPI‑prod.
We further evaluate our framework against DIAL‑MPC (Xue et al., 2025), an annealed MPPI variant that adapts its covariance using diffusion‑inspired updates (Ho et al., 2020; Chi et al., 2024). Although DIAL-MPC is significantly more efficient than vanilla MPPI, it still relies on rejection sampling to eliminate infeasible trajectories—because its experts are fused at the function level and have no direct awareness of the feasible solution space. By contrast, our solution-level fusion lets us steer the sampler toward valid regions before any draws are discarded. To highlight this advantage, we embed DIAL-MPC within our TT-PoE pipeline (denoted TT-PoE-DIAL) and show that this hybrid consistently outperforms both standalone DIAL-MPC and our original TT-PoE-MPPI in terms of sampling efficiency and overall task success.
Table 5 reports success rate, total cost, and steps-to-goal for each method. Integrating TT with both MPPI and DIAL‑MPC increases success rates while reducing cost and required steps. Notably, the TT‑PoE‑MPPI‑prod variant matches or exceeds the performance of TT‑PoE‑MPPI, validating the effectiveness of learning and summing individual feasibility distributions within a product‑of‑experts framework. We define task success as reaching the target in fewer than 25 steps with total cost below $10^{6}$. Further experimental details are provided in the Appendix.
@@ 5.5 Higher Dimensions and Alternative Optimality Distributions
Table 5: Comparison of different methodologies in whole body obstacle avoidance task across 100 random obstacle locations and initial and target configurations using 1024 sample size. Values are normalized to the MPPI method. Thus, steps and cost values for MPPI are zero. Negative values indicate superior performance, with bold formatting highlighting the best results.
Method succ. rate log(# $\text{steps}_{n}$) log($\text{cost}_{n}$)
MPPI $30\%$ $0.00$ $0.00$
$T$-PoE-MPPI $53\%$ 0.10-0.10 0.09-0.09
$T$-PoE-MPPI-prod $50\%$ 0.14-0.14 0.11-0.11
DIAL $88\%$ 0.35-0.35 0.36-0.36
$T$-PoE-DIAL $\bm{89\%}$ $\bm{-0.50}$ $\bm{-0.53}$